1. Find p such that P( X > 2 ) = 1/2 where X is Geometric( p )
Answer #1

Given X→ GCP) To ind oa P suck that e lcoous thot 2 + 93 = 1,2 (a)s Cl/2)V3 0.7937 = 1-9 0.20 63